OpenStudy (aaronandyson):

5 cos (A) = 6 sin (A) Find tan(A)

OpenStudy (aaronandyson):

also find

OpenStudy (aaronandyson):

\[\frac{ 12 \sin (A) - 3 \cos(A) }{ 12 \sin (A) + 3 \cos (A) }\]

OpenStudy (anonymous):

\[\tan (A) = \frac{ \sin (A) }{ \cos (A) }\] \[\tan (A) = \frac{ 5 }{ 6 }\]

OpenStudy (aaronandyson):

I got 6/5?

OpenStudy (anonymous):

how 6/5?

OpenStudy (aaronandyson):

Sorry I had made a mistake yes now what??

OpenStudy (aaronandyson):

@iambatman @kiwisallday

OpenStudy (anonymous):

the second part is a continuation of the first question?

OpenStudy (aaronandyson):

yes.

OpenStudy (anonymous):

\[\tan (A) = opposite/adjacent = 5/6\]

OpenStudy (anonymous):